# Fine-tuning with Very Large Dropout

### v2

@@ Abstract
It is impossible today to pretend that the practice of machine learning is compatible with the idea that training and testing data follow the same distribution. Several authors have recently used ensemble techniques to show how scenarios involving multiple data distributions are best served by representations that are both richer than those obtained by regularizing for the best in-distribution performance, and richer than those obtained under the influence of the implicit sparsity bias of common stochastic gradient procedures.
@@ 1 Introduction
The dropout technique (Srivastava et al., 2014) seems well suited to finding weakly relevant features because randomly masking units of a representation layer during training amounts to forming random subsets of all other available features. However, in order to form small subsets, one would have to use very high levels of dropout. Unfortunately, training a sizable deep network from scratch with such a large dropout is practically impossible. Instead, computationally demanding methods, such as adversarial sampling (Zhang et al., 2022; Chen et al., 2023) and representation ensembles (Zhang & Bottou, 2023), have been proposed to find weakly relevant features while training a network from scratch.
@@
In the present work, we adopt the out-of-distribution fine-tuning setup (three-distributions) of Ramé et al. (2022b). In this framework, we have access to a model pre-trained using a large dataset for a task weakly related to the task of interest. This pre-trained model is then fine-tuned on datasets that illustrate the task of interest, and then tested on a dataset for the same task but with a different distribution. However, instead of enriching the representations by constructing ensembles (Zhang & Bottou, 2023) or averaging weights (Ramé et al., 2022b; a; Wortsman et al., 2022b), we simply fine-tune using very large dropout levels, randomly masking above 90% of the units in the representation layer. We find that this simple approach exceeds the performance of both ensemble and weight-averaging methods. This result is not only practically meaningful, but also clarifies the idea of richer representation.
@@ 2 Related Work
Researchers have devised ways to obtain more versatile representations by engineering a diversity of datasets, architectures, and even hyper-parameters (Chen et al., 2020; Wang et al., 2022; Dvornik et al., 2020; Bilen & Vedaldi, 2017; Gontijo-Lopes et al., 2021; Li et al., 2021; 2022; Chowdhury et al., 2021), as an alternative to the most popular approach which consists of simply using ever larger datasets (Bommasani et al., 2021b).
@@
Although modern deep residual networks feature highly complex nonconvex cost functions, several authors have shown that their final training phase remains mostly confined to a nearly-convex attraction basin (Izmailov et al., 2018; Li et al., 2018; Frankle et al., 2020). The same observation holds when fine-tuning a large pre-trained network using a dataset whose size is considerably smaller than the dataset size one would need to train such a large network from scratch. As long as one starts from the same pre-trained model, Wortsman et al. (2022a) and Ramé et al. (2022b; a) observe that averaging the weights of diverse fine-tuned models can reproduce the i.i.d. and o.o.d. performances of the ensemble of these models, implying that fine-tuning is a near-linear process.
Maddox et al. (2021) and Mu et al. (2019) propose instead to approximate the fine-tuning process with a first-order Taylor expansion, obtaining a linear system operating on top of the ntk features. Evci et al. (2022) match the performance of fine-tuning by merely learning a strongly regularized linear model that takes all internal layer states as inputs. Meanwhile, Yu et al. (2023) efficiently fine-tune large foundational language models by essentially restricting the weight updates to low dimensional manifolds.
@@ 3.1 The three-distributions setup
The two-distributions setup is a commonly used for transfer learning. In this setup, features $\Psi$ are obtained by pre-training a network on a large training set associated with a first distribution $\mathcal{T}_{\text{p}}$. These features are then used to construct or initialize a new model $\omega_{\text{d}}\circ\Psi$, which is then trained using a smaller training set associated with a second distribution $\mathcal{T}_{\text{d}}$. The question is to determine which pre-training approach is most likely to make the features $\Psi$ useful for the transfer task $\mathcal{T}_{\text{d}}$.
@@ 3.2 Method
Fine-tuning is carried out with a standard stochastic learning procedure (e.g. sgd or adam) after applying a very large dropout to the penultimate layer representation $\Phi$. Unlike (Srivastava et al., 2014), we only apply dropout on the penultimate layer representation $\Phi$, [^1] Except in Figure 6 in the appendix where we investigate the effect of dropping entire inner residual blocks during fine-tuning. because skip connections in residual networks expose many inner-layer features to the last linear layer, as illustrated by the decomposition of residual networks proposed by Veit et al. (2016),
@@
where $\odot$ represents the component-wise product and $m(\lambda)$ is a vector of random Bernoulli variables equal to 00 with probability $\lambda$ and $1$ with probability $1-\lambda$. The additive decomposition of $\Phi(x)$ in equation equation 1 makes clear that applying dropout to $\Phi(x)$ simultaneously blocks the contributions $\phi_{i}(x)$ of all residual blocks.
@@ 4 Experiments
We perform most experiments using pacs (Li et al., 2017), vlcs (Fang et al., 2013), office home (Venkateswara et al., 2017), and terra  incognita (Beery et al., 2018) datasets. These datasets are part of the DomainBed suite which has been widely used for this kind of experiments. [^2] https://github.com/facebookresearch/DomainBed With $9,991$ to $24,788$ examples, they are substantially smaller than the pre-training dataset ImageNet with 1.2m examples. We also use the larger domainnet dataset (Peng et al., 2019), 0.58m examples, to show that linear connectivity breaks down when the fine-tuning dataset size becomes comparable with the pre-training dataset size and justifies carrying out many more fine-tuning iterations.
@@
Fine-tuning experiments are carried out on two widely used residual architectures: convolutional residual networks, ResNets (He et al., 2016), and vision transformers, ViTs (Dosovitskiy et al., 2020). Unless otherwise stated, all convolutional residual network experiments are carried out using the ResNet50 neural network pre-trained on ImageNet with substantial data augmentations [^3] https://pytorch.org/blog/how-to-train-state-of-the-art-models-using-torchvision-latest-primitives/ such as trivialaugment (Müller & Hutter, 2021), cutmix (Yun et al., 2019), and random erasings (Zhong et al., 2020). These augmentations mimic the properties of large foundational models that learn substantial diverse features using very large and diverse pre-training data. We refer to this network as ResNet50 #2 as opposed to the original ResNet50 #1 recipe described by He et al. (2016). Similarly, all visual transformer fine-tuning experiments leverage a ImageNet pre-trained ViT-L-16 model [^4] https://github.com/pytorch/vision/tree/main/references/classification#vit_l_16 with 304m parameters.
Figure 1: o.o.d. performance comparison between very large dropout, ensembles, and weight averaging methods on four DomainBed tasks and two backbones (ResNet50 and ViT-L-16). Baseline results were obtained using plain fine-tuning with different hyperparameters. Weight averaging results either average the model weights collected every 300 iterations along each fine-tuning trajectory or the final model weights of all fine-tuning trajectories as in (Ramé et al., 2022b). Ensemble results average instead the model outputs. Finally, large dropout results were obtained like the baseline results but using a 90% dropout rate on the penultimate layer. Each box summarizes the results obtained with 1296 hyper-parameters combinations (no hyper-parameter selection).
@@
dataset baseline weight avg (single) ensemble (single) large dropout weight avg (multi) ensemble (multi)
@@
4.1 Very large dropout yields better o.o.d. performance
Using these same datasets, Gulrajani & Lopez-Paz (2020) argue that plain Empirical Risk Minimization (erm) equals and often betters the o.o.d. performance of purposefully designed methods, such as coral (Sun & Saenko, 2016), vrex (Krueger et al., 2021), and irm (Arjovsky et al., 2019). More recently, Arpit et al. (2022), Cha et al. (2021), Ramé et al. (2022b), and Ramé et al. (2022a) find that ensemble and weight averaging methods consistently outperform the o.o.d. performance of erm. We now show that fine-tuning with very large dropout outperforms the o.o.d. performance of these state-of-the-art methods. [^5] Code: https://github.com/TjuJianyu/verylarge_dropout
Following these earlier works, we focus on the o.o.d. performance of these methods because this is the testing performance that matters for the practical situations that the three-distribution setup targets. The i.i.d. performance must then be viewed as a secondary indicator, akin to the training set performance, which could be used to cast a light on the training process but cannot be used as a predictor of the system performance under its actual testing distribution.
Figure 1 reports the o.o.d. performances of the following approaches.
  • Baseline results are obtained by fine-tuning our ResNet50 or ViT-L-16 using sgd with 0.9 momentum for $10,000$ iterations. [^6] We use a batch size $32$ for all ResNet fine-tunings, and reduce the batch size to $16$ for all ViT-L-16 fine-tunings due to the vram constraint. A 10% learning rate decay is applied at $5000^{th}$ iterations. For each choice of three training sub-datasets, we repeat three experiments for each combination of learning rate in $\{10^{-3},5.10^{-4}\}$ and L2 weight decay in $\{10^{-4},5.10^{-5},10^{-5}\}$. We measure the i.i.d. performance using 20% examples held from the training data, and we measure o.o.d. performance on the fourth sub-dataset. Each box summarizes the performances obtained using two learning rate choices and three weight decay choices, for four possible training dataset choices, resulting in $(2\times 3)^{4}=1296$ results. Following Gulrajani & Lopez-Paz (2020), we prevent overfitting by early-stopping at the best i.i.d. performance.
  • Dropout results are obtained using the same protocol but using a 90% dropout rate on the penultimate layer representation.
  • Ensemble results are obtained in two ways, either using an ensemble of checkpoints collected along each fine-tuning trajectory, or using the ensemble of the final checkpoints collected along all fine-tuning trajectories with different hyper-parameters.
  • Weight averaging results approximate the corresponding ensembling results by averaging the model weights instead of averaging the model outputs.
As expected, both ensemble methods (Ueda & Nakano, 1996; Dietterich, 2000) and their weight averaging approximation (Ramé et al., 2022b; Wortsman et al., 2022a) improve on the o.o.d. baseline performance. However, fine-tuning with a very large dropout outperforms the o.o.d. performance of both ensemble and weight averaging methods. There is even a large gap between the worst dropout results and the best ensemble results for the office home and pacs datasets. Figure 7 in the Appendix shows that the i.i.d. performance of the large dropout method lags behind that of ensembles, revealing that the o.o.d. performance improvement is not a secondary effect of some i.i.d. performance improvement (i.e. the o.o.d. performance gaps in o.o.d. Figure 1 do not come from i.i.d. overfitting/underfitting.).
The box plots in Figure 1 summarize the o.o.d. performance of all possible hyper-parameter combinations. Following again Gulrajani & Lopez-Paz (2020), Table 1 features o.o.d. performances obtained using the hyper-parameter combination that provides the best i.i.d. performance. Because ResNet50 produces a better performance than ViT-L-16 on these o.o.d. fine-tuning tasks, our experiments in the following sections will be conducted on ResNet50.
4.1.1 Optimal o.o.d. dropout rate
To the best of our knowledge, such large dropout rates (90% and above) are considered unsuitable for training a network from scratch and have not been previously used for fine-tuning either. This section illustrates how the optimal dropout rate can be very high in fine-tuning scenarios and falls to small values when one gets closer to training the network from scratch.
Figure 2 compares various dropout rates on the four DomainBed tasks. A $90\%$ dropout rate reliably produces good o.o.d. performance on all four tasks. The optimal dropout rate for o.o.d. performance ranges from 90% to 95% for vlcs and pacs task (with 10k examples). And becomes slightly smaller, about 90%, for the slighlty larger datasets office home and terra incognita (with 15k to 25k examples).
Increasing the fine-tuning dataset size to approach the pre-training dataset size
The larger the fine-tuning dataset, the more fine-tuning iterations we can make without overfitting. When the fine-tuning dataset size approaches the pre-training dataset size, the difference between fine-tuning and training from scratch becomes less clear, the linear connectivity property disappears, the linear approximation perspective on fine-tuning no longer holds, and the optimal dropout rate falls sharply. Figure 4 illustrates this effect using the larger domainnet dataset (Peng et al., 2019) that contains 586k examples (almost half as big as ImageNet) and requires $30,000$ fine-tuning iterations.
Training from scratch on the fine-tuning dataset
Figure 4 shows the effect of various dropout rates when one trains a network on the vlcs task from scratch, that is starting from a randomly initialized network without pretraining (i.e. the pretraining dataset size is zero). The optimal dropout rate falls to about zero. Dropout rates higher than 50% have a negative impact on both the i.i.d. and the o.o.d. performance of the network. This suggests again that high dropout rates make it difficult to create new features (a nonlinear operation), but does not prevent leveraging existing features that were possibly buried in the network inner layers (a linear operation). This is the idea of richer representation we discussed in section 1. We will provide other fine-tuning experiments on the richer representation idea in Section 4.3.
Figure 2: Effect of diverse dropout rates during fine-tuning. The best o.o.d. performances are attained using rates around or above 90%. A large dropout rate (e.g. 90%) reliably produces good o.o.d. performance on all four tasks.
Figure 3: Comparison of various dropout rates on the larger domainnet dataset (586K examples), whose size approaches the pretraining dataset size (ImageNet, 1.2M examples). The optimal dropout rate falls to about 50%, a value comparable to the dropout rates traditionally used when training from scratch.
4.2 Popular fine-tuning techniques do not substantially improve the o.o.d. performance of large dropouts
Various fine-tuning techniques have been proposed to improve the fine-tuning ability to leverage the representations learned by a pre-trained model, such as using a larger learning rate on the last layer (Caron et al., 2020; Bardes et al., 2021; Kumar et al., 2022) or, as discussed above, using weight averaging and ensemble methods (Ramé et al., 2022b; a; Arpit et al., 2022). We show in this section that using these techniques in addition to very large dropout rates do not yield much o.o.d. performance improvements over using large dropout rates alone. Note that we still expect some incremental benefits because both weight averaging and ensembles reduce the stochastic optimization noise and accelerate training in general (Polyak & Juditsky, 1992).
4.2.1 Large learning rates for the last layer
dataset baseline baseline + 10$\times$ last-layer lr baseline + large dropout baseline + large dropout + 10$\times$ last-layer lr

### v3

# Finetuning with Very-large Dropout

@@ Abstract
It is impossible today to pretend that the practice of machine learning is always compatible with the idea that training and testing data follow the same distribution. Several authors have recently used ensemble techniques to show how scenarios involving multiple data distributions are best served by representations that are both richer than those obtained by regularizing for the best in-distribution performance, and richer than those obtained under the influence of the implicit sparsity bias of common stochastic gradient procedures.
@@ 1 Introduction
The dropout technique (Srivastava et al., 2014) seems well suited to find weakly relevant features because randomly masking units of a representation layer during training amounts to forming random subsets of all other available features. However, in order to form small subsets, one would have to use very high levels of dropout. Unfortunately, training a sizable deep network from scratch with such a large dropout is practically impossible. Instead, computationally demanding methods, such as adversarial sampling (Zhang et al., 2022; Chen et al., 2023) and representation ensembles (Zhang & Bottou, 2023), have been proposed to find weakly relevant features while training a network from scratch.
@@
In the present work, we adopt the out-of-distribution fine-tuning setup (three-distributions) of Ramé et al. (2022b). In this framework, we have access to a model pre-trained using a large dataset for a task weakly related to the task of interest. This pre-trained model is then fine-tuned on datasets that illustrate the task of interest, and then tested on a dataset for the same task but with a different distribution. However, instead of enriching the representations by constructing ensembles (Zhang & Bottou, 2023) or averaging weights (Ramé et al., 2022b, a; Wortsman et al., 2022b), we simply fine-tune using very large dropout levels, randomly masking above 90% of the units in the representation layer. We find that this simple approach exceeds the performance of both ensemble and weight-averaging methods. This result is not only practically meaningful, but also clarifies the idea of richer representation.
@@ 2 Related Work
Researchers have devised ways to obtain more versatile representations by engineering a diversity of datasets, architectures, and even hyper-parameters (Chen et al., 2020; Wang et al., 2022; Dvornik et al., 2020; Bilen & Vedaldi, 2017; Gontijo-Lopes et al., 2021; Li et al., 2021, 2022; Chowdhury et al., 2021), as an alternative to the most popular approach which consists of simply using ever larger datasets (Bommasani et al., 2021b).
@@
Although modern deep residual networks feature highly complex nonconvex cost functions, several authors have shown that their final training phase remains mostly confined to a nearly-convex attraction basin (Izmailov et al., 2018; Li et al., 2018c; Frankle et al., 2020). The same observation holds when fine-tuning a large pre-trained network using a dataset whose size is considerably smaller than the dataset size one would need to train such a large network from scratch. As long as one starts from the same pre-trained model, Wortsman et al. (2022a) and Ramé et al. (2022b, a) observe that averaging the weights of diverse fine-tuned models can reproduce the i.i.d. and o.o.d. performances of the ensemble of these models, implying that fine-tuning is a near-linear process.
Maddox et al. (2021) and Mu et al. (2019) propose instead to approximate the fine-tuning process with a first-order Taylor expansion, obtaining a linear system operating on top of the ntk features. Evci et al. (2022) match the performance of fine-tuning by merely learning a strongly regularized linear model that takes all internal layer states as inputs. Meanwhile, (Yu et al., 2023) efficiently fine-tune large foundational language models by essentially restricting the weight updates to low dimensional manifolds.
@@ 3.1 The three-distributions setup
The two-distributions setup is commonly used for transfer learning. In this setup, features $\Psi$ are obtained by pre-training a network on a large training set associated with a first distribution $\mathcal{T}_{\text{p}}$. These features are then used to construct or initialize a new model $\omega_{\text{d}}\circ\Psi$, which is then trained using a smaller training set associated with a second distribution $\mathcal{T}_{\text{d}}$. The question is to determine which pre-training approach is most likely to make the features $\Psi$ useful for the transfer task $\mathcal{T}_{\text{d}}$.
@@
3.2 Examples
Considering a logistic regression with parameter $\omega\in\mathbb{R}^{n}$ operating on a vector $\Psi$ of $n$ features and predicting a binary target $Y$ representing our ${\mathcal{T}}_{\text{d}}$ distribution. Assume further that each individual feature $\Psi_{i},i\in[1,\dots,n]$ perfectly predicts $Y$, that is, zero classification error can be achieved with a regression $\omega$ whose only nonzero parameter marks the $i$-th feature. During gradient-based optimization, achieving zero loss by using only one feature prevents the system from using the other features, because of the “gradient starving” phenomenon (Pezeshki et al., 2021). We now evaluate this trained system on a target distribution $\tilde{\mathcal{T}}_{\text{d}}$ that only differs from ${\mathcal{T}}_{\text{d}}$ because some features were missing and have been replaced by zeroes. If our trained system (on ${\mathcal{T}}_{\text{d}}$) depends only on one feature, we better hope that this is not one of the missing ones in target distribution $\tilde{\mathcal{T}}_{\text{d}}$.
In this linear case, the following three strategies are equivalent in terms of encouraging the optimization process to learn more features: 1) feature-bagging (ensemble) (Bryll et al., 2003); 2) Dropout; 3) $L_{2}$ regularization on $\omega$ (Check Srivastava et al. (2014) for the proof). We know that the feature-bagging approach solves the problem above by construction. Thus, in the linear case, all three strategies solve the above problem.
In the case of a multilayer network, however, this equivalence is broken. In particular, $L_{2}$ regularization on the inner layer parameters plays the different role of encouraging sparse representations (Blanc et al., 2020; Andriushchenko et al., 2023). Dropout and deep ensembles may achieve comparable error rates in distribution but differ sharply when it comes to estimating prediction uncertainty (Ashukha et al., 2020). These differences become very important when one fine-tunes the model using out-of-distribution data, making deep ensembles and weight averaging ensembles more attractive than dropout for o.o.d. generalization (Ramé et al., 2022b, a; Wortsman et al., 2022a; Cha et al., 2021; Arpit et al., 2022).
Our contribution shows that using a very large dropout rate during fine-tuning (rather than during initial training) substantially improves on the o.o.d. performance of both ensemble and weight-averaging. This simple approach was not considered before, possibly because such large dropout rates are not usable during pre-training, resulting in poor performance overall.
@@ 3.3 Method
Fine-tuning is carried out with a standard stochastic learning procedure (e.g. sgd or adam) after applying a very large dropout to the penultimate layer representation $\Phi$. Unlike (Srivastava et al., 2014), we only apply dropout on the penultimate layer representation $\Phi$, because skip connections in residual networks expose many inner-layer features to the last linear layer, as illustrated by the decomposition of residual networks proposed by Veit et al. (2016),
@@
where $\odot$ represents the component-wise product and $m(\lambda)$ is a vector of random Bernoulli variables equal to 00 with probability $\lambda$ and $1$ with probability $1-\lambda$. The additive decomposition of $\Phi(x)$ in equation (3.3) makes clear that applying dropout to $\Phi(x)$ simultaneously blocks the contributions $\phi_{i}(x)$ of all residual blocks.
In this work, this approach is called very-large dropout, because the dropout rate ($\sim$90%) is far larger than people used before.
@@ 4 Experiments
Dataset
We perform most experiments using pacs (Li et al., 2017), vlcs (Fang et al., 2013), office home (Venkateswara et al., 2017), and terra  incognita (Beery et al., 2018) datasets. These datasets spam in diverse domains, from wild images with different environment conditions to artificial sketching and painting, from natural animals to home furniture. With $9,991$ to $24,788$ examples, these datasets are substantially smaller than the pre-training dataset ImageNet with 1.2m examples.
@@
Models
We carry out experiments using two wisely used residual architectures. For the convolutional network experiments, we use a ResNet50 architecture (He et al., 2016) with 25m parameters. [^1] https://pytorch.org/blog/how-to-train-state-of-the-art-models-using-torchvision-latest-primitives/ For the visual transformer experiments, we use the large vision transformer ViT-L-16 (Dosovitskiy et al., 2020) with 304m parameters. [^2] https://github.com/pytorch/vision/tree/main/references/classification#vit_l_16
Pre-training
Unless otherwise stated, all experiments are carried out using networks pre-trained using refined data augmentations initially introduced in the context of residual networks: trivialaugment (Müller & Hutter, 2021), cutmix (Yun et al., 2019), and random erasings (Zhong et al., 2020). We use these augmentations to mimic the properties of large foundational models trained using very large and diverse pre-training data.
Baselines
Using these same datasets, Gulrajani & Lopez-Paz (2020) argue that plain Empirical Risk Minimization (erm) equals and often betters the o.o.d. performance of purposefully designed methods, such as coral (Sun & Saenko, 2016), dro (Sagawa et al., 2019), mldg (Li et al., 2018a), dann (Ganin et al., 2015), c-dann (Li et al., 2018d), mmd (Li et al., 2018b), vrex (Krueger et al., 2021), and irm (Arjovsky et al., 2019). More recently, Arpit et al. (2022), Cha et al. (2021), Ramé et al. (2022b), and Ramé et al. (2022a) find that ensemble and weight averaging methods consistently outperform the o.o.d. performance of erm.
Therefore, it is sufficient to compare our results with those of the ensemble, weight averaging, and erm methods which are the strongest available baselines. [^3] Gulrajani & Lopez-Paz (2020); Arpit et al. (2022); Cha et al. (2021); Ramé et al. (2022b, a) provide the details about how ensemble and weighting averaging outperform other baseline methods.
@@
dataset erm weight average (single run) ensemble (single run) very-large dropout weight average (multi run) ensemble (multi run)
@@
[…]
dataset erm 10$\times$ last-layer lr very-large dropout very-large dropout + 10$\times$ last-layer lr
@@
4.1 Very large dropout yields better o.o.d. performance
Table 1 shows our main results that comparing our very-large dropout approach and baseline methods on four o.o.d. datasets and two pretrained backbones.  [^4] Code: https://github.com/TjuJianyu/verylarge_dropout
erm
results are obtained by fine-tuning ResNet50 or ViT-L-16 using sgd with 0.9 momentum for $10,000$ iterations. [^5] We use a batch size $32$ for all ResNet fine-tunings, and reduce the batch size to $16$ for all ViT-L-16 fine-tunings due to the vram constraint. A 10% learning rate decay is applied at $5000^{th}$ iterations. For each choice of three training sub-datasets, we repeat three experiments for each combination of learning rate in $\{10^{-3},5.10^{-4}\}$ and L2 weight decay in $\{10^{-4},5.10^{-5},10^{-5}\}$. Following Gulrajani & Lopez-Paz (2020), we prevent overfitting by early-stopping on 20% hold-out i.i.d. validation examples, select hyperparameter (for each choice of training sub-datasets) according to the best i.i.d. performance. Finally, we evaluate the selected models on the fourth sub-dataset and average the four choices of training sub-datasets.
Ensemble (single run)
results are obtained by an ensemble of checkpoints collected (every 300 iterations) along each fine-tuning trajectory.
Weight average (single run)
results approximate the corresponding ensemble (single run) results by averaging the model weights instead of averaging the model outputs.
Ensemble (multi run)
results are obtained by an ensemble of final checkpoints collected along all fine-tuning trajectories with different hyper-parameters ($2\times 3=6$ in total).
Weight average (multi run)
results approximate the corresponding ensemble (multi run) results by averaging the model weights.
Very-large dropout
results are obtained using the same protocol but using a 90% dropout rate on the penultimate layer representation.
As expected, both ensemble methods (Ueda & Nakano, 1996; Dietterich, 2000) and their weight averaging approximation (Ramé et al., 2022b; Wortsman et al., 2022a) improve on the o.o.d. erm performance. However, fine-tuning with a very large dropout outperforms the o.o.d. performance of both ensemble and weight averaging methods.
Because ResNet50 produces a better performance than ViT-L-16 on these o.o.d. fine-tuning tasks, our experiments in the following sections will be conducted on ResNet50.
4.2 Very-large dropout $+$ other fine-tuning techniques
Various fine-tuning techniques have been proposed to improve the fine-tuning ability to leverage the representations learned by a pre-trained model, such as using a larger learning rate on the last layer (Caron et al., 2020; Bardes et al., 2021; Kumar et al., 2022) or, as discussed above, using weight averaging and ensemble methods (Ramé et al., 2022b, a; Arpit et al., 2022). In this section, we show that incorporating these techniques in additional to very-large dropout can further enhance o.o.d. performance, i.e. very-large dropout approach is compatible to these existing fine-tuning techniques.
More importantly, very-large dropout approach dominates the o.o.d. performance improvements. i.e., all these finetuning techniques do not yield much o.o.d. performance improvements over using large dropout rates alone.
Table 3: Very-large dropout $+$ ensembles or weight averagings. The erm and very-large dropout results are the same as those reported in Table 1. In contrast, the ensemble and weight averaging results are now obtained by averaging the output or the weights of models fine-tuned with large dropouts. Ensemble and weight averaging techniques provide a marginal o.o.d. performance improvement on vlcs or office home and a negligible o.o.d. performance improvement on pacs or terra incognita.
dataset erm very-large dropout very-large dropout + weight average (single run) very-large dropout + ensemble (single run) very-large dropout + weight average (multi run) very-large dropout + ensemble (multi run)
vlcs 78.3 80.1 80.6 80.5 80.4 80.3
office home 71.4 73.6 74.2 74.3 74.4 74.2
pacs 87.3 88.5 88.6 88.8 89.0 89.0
terra incognita 51.0 53.9 54.0 54.7 52.3 54.7
Average 72.0 74.0 74.4 74.6 74.0 74.6
4.2.1 Very-large dropout
$+$ large learning rates for the last layer